Using the same sample of gas ( P 1 = 795 torr , T 1 = 25 ∘ C ), we wish to change the pressure to 7950 torr with no accompanying change in volume or amount of gas. What temperature T 2 , in Celsius, is needed to reach this pressure?
Answer #1

Given:

Ti = 25.0 oC

= (25.0+273) K

= 298 K

Pi = 795 torr

Pf = 7950 torr

use:

Pi/Ti = Pf/Tf

795.0 torr / 298.0 K = 7950.0 torr / Tf

Tf = 2980 K

= (2980 - 273) oC

= 2707 oC

Answer: 2707 oC